# Hierarchy of quantum correlations in qubit-qutrit axially symmetric states

### v1

@@ Abstract
We investigate quantum correlations in a hybrid qubit–qutrit system subject to both axial and planar single-ion anisotropies, dipolar spin–spin interactions, and Dzyaloshinskii–Moriya (DM) coupling. Using Negativity, Measurement-Induced Nonlocality (MIN), Uncertainty-Induced Nonlocality (UIN), and Bell nonlocality (as quantified by the CHSH inequality) as measures, we analyze the interplay between anisotropy parameters, magnetic fields, and temperature on the survival of quantum correlations. Our results demonstrate that Bell nonlocality and entanglement (Negativity) are highly sensitive to temperature and anisotropy, exhibiting sudden death under thermal noise, whereas MIN and UIN are significantly more robust. In particular, these discord-like and information-theoretic measures provide the largest baseline and persists even in parameter regions where entanglement vanishes, highlighting its suitability as a quantumness witness in realistic conditions. Notably, our Bell nonlocality study is tailored to the asymmetric qubit–qutrit setting by exploiting a recently developed qubit–qudit CHSH maximization framework. However, Bell nonlocality is confirmed to be the most fragile, surviving only in narrow parameter windows at low temperature. A key finding of this work is that we observe the fragility hierarchy
$\color[rgb]{0,0,0}\text{Bell nonlocality}<\text{Negativity}<\text{UIN}(\text{MIN})$}
@@ 4 Results and Discussions
At $T=0$, the state is maximally entangled, yielding the highest Negativity. As $T$ increases, thermally induced mixing reduces the off‑diagonal coherences responsible for entanglement, causing a smooth, monotonic decay of Negativity. At finite temperature, “entanglement‑death” is reached when Negativity falls to zero. MIN also decreases with temperature, but it persists beyond entanglement‑death, reflecting that certain nonclassical correlations survive even in separable states. Its decay is gentler than that of Negativity, vanishing significantly only at higher temperatures. Defined via the Wigner–Yanase skew information, UIN begins at a larger value than MIN at $T=0$. Its thermal decay parallels that of MIN and Negativity, but does not exhibit crossover behavior between $J_{z}=1,2,3$. Specifically, $J_{z}=1$ has the highest correlations at $T=0$ but decays first. For the same set of parameters, this $J_{z}=1$ decay mirrors the behavior of LQU and LQFI in Ref. [47], which likewise decrease without any sudden jumps as $T$ rises. A key difference is that Negativity and Bell–CHSH reach zero and cross the classical limit 2, respectively, at lower temperature, whereas LQU/LQFI (discord-like measures) approach zero more gradually. The CHSH parameter attains its maximum only in the very low‑temperature regime. Even moderate thermal fluctuations reduce the Bell–CHSH value below the classical bound of 2, such that Bell nonlocality is lost. This hierarchy of fragility — $\text{Bell nonlocality}<\text{Negativity}<\text{UIN}(\text{MIN})$ — highlights the increasing robustness of more general nonclassical correlations against thermal noise. Furthermore, the hierarchy suggests that the UIN (MIN) is a better resource against thermal fluctuations in a thermally activated system.
@@
Further, in Fig. 2 (a–d), we study the effect of increasing temperature while varying the isotropic exchange coupling $J$ from the antiferromagnetic to the ferromagnetic regime, while keeping $J_{z}=1$ and all other parameters fixed. For $J=0$, the spins are uncoupled beyond the $s_{z}S_{z}$ term. As $J$ becomes increasingly negative (ferromagnetic $XY$ coupling), negativity increases, reflecting the stronger singlet-like correlations favored by the transverse exchange. In this case as well, entanglement death is observed with rising temperature, indicating that Negativity is highly susceptible to thermal fluctuations. MIN follows a qualitatively similar trend but remains nonzero even for smaller values of $J$ at higher $T$, underscoring its sensitivity to nonclassical disturbances beyond strict entanglement. The high $T$ “tail” of MIN is notably broader. At $T=0$, UIN again starts from a higher baseline and shows no abrupt thresholds. Its monotonic behavior across the entire $T$-range demonstrates its robustness as a measure of quantum correlations derived from information-theoretic uncertainty. Even at low temperature, a minimum ferromagnetic strength of J=0.7J=-0.7 is not sufficient for the CHSH parameter to exceed the classical limit. With the variation of $J$, the fragility of the quantumness measures follows the same hierarchy as in the previous case, namely $\text{Bell nonlocality}<\text{Negativity}<\text{UIN}(\text{MIN})$. This demonstrates that Bell nonlocality is the most challenging to sustain, while MIN and UIN serve as more reliable indicators of nonclassical correlations even at finite temperatures.
@@
Comparing Figs. 3 and 4 shows that for the strong ferromagnetic transverse coupling (J=2.5J=-2.5) and negative longitudinal exchange ($J_{z}=-1$), tuning the qubit Zeeman term ($B_{1}$) produces larger low-temperature entanglement and a wider Bell-violation window than tuning the qutrit Zeeman term ($B_{2}$). In both scans, MIN and UIN remain appreciable at higher temperatures, where entanglement and Bell nonlocality have already been suppressed, with UIN being the most thermally robust information-theoretic signature. The qualitative fragility ordering Bell << Negativity << UIN (MIN) remains valid across both parameter sweeps. However, the exact field windows and temperature thresholds differ depending on whether the qubit or qutrit field is being varied.
@@ 5 Conclusions and outlook
$\text{Bell nonlocality}<\text{Negativity}<\text{UIN}(\text{MIN}),$

### v2

@@ Abstract
We investigate quantum correlations in a hybrid qubit–qutrit system subject to both axial and planar single-ion anisotropies, dipolar spin–spin interactions, and Dzyaloshinskii–Moriya (DM) coupling. Using Negativity, Measurement-Induced Nonlocality (MIN), Uncertainty-Induced Nonlocality (UIN), and Bell nonlocality (as quantified by the CHSH inequality) as measures, we analyze the interplay between anisotropy parameters, magnetic fields, and temperature on the survival of quantum correlations. Our results demonstrate that Bell nonlocality and entanglement (Negativity) are highly sensitive to temperature and anisotropy, exhibiting sudden death under thermal noise, whereas MIN and UIN are significantly more robust. In particular, these discord-like and information-theoretic measures provide the largest baseline and persist even in parameter regions where entanglement vanishes, highlighting its suitability as a quantumness witness in realistic conditions. Notably, our Bell nonlocality study is tailored to the asymmetric qubit–qutrit setting by exploiting a recently developed qubit–qudit CHSH maximization framework. However, Bell nonlocality is confirmed to be the most fragile, surviving only in narrow parameter windows at low temperature. A key finding of this work is that we observe the fragility hierarchy
$\color[rgb]{0,0,0}\text{Bell nonlocality}\subseteq\text{Negativity}\subseteq\text{UIN}(\text{MIN})$}
@@ 4 Results and Discussions
At $T=0$, the state is maximally entangled, yielding the highest Negativity. As $T$ increases, thermally induced mixing reduces the off‑diagonal coherences responsible for entanglement, causing a smooth, monotonic decay of Negativity. At finite temperature, “entanglement‑death” is reached when Negativity falls to zero. MIN also decreases with temperature, but it persists beyond entanglement‑death, reflecting that certain nonclassical correlations survive even in separable states. Its decay is gentler than that of Negativity, vanishing significantly only at higher temperatures. Defined via the Wigner–Yanase skew information, UIN begins at a larger value than MIN at $T=0$. Its thermal decay parallels that of MIN and Negativity, but does not exhibit crossover behavior between $J_{z}=1,2,3$. Specifically, $J_{z}=1$ has the highest correlations at $T=0$ but decays first. For the same set of parameters, this $J_{z}=1$ decay mirrors the behavior of LQU and LQFI in Ref. [47], which likewise decrease without any sudden jumps as $T$ rises. A key difference is that Negativity and Bell–CHSH reach zero and cross the classical limit 2, respectively, at lower temperature, whereas LQU/LQFI (discord-like measures) approach zero more gradually. The CHSH parameter attains its maximum only in the very low‑temperature regime. Even moderate thermal fluctuations reduce the Bell–CHSH value below the classical bound of 2, such that Bell nonlocality is lost. This hierarchy of fragility — $\text{Bell nonlocality}\subseteq\text{Negativity}\subseteq\text{UIN}(\text{MIN})$ — highlights the increasing robustness of more general nonclassical correlations against thermal noise. Furthermore, the hierarchy suggests that the UIN (MIN) is a better resource against thermal fluctuations in a thermally activated system.
@@
Further, in Fig. 2 (a–d), we study the effect of increasing temperature while varying the isotropic exchange coupling $J$ from the antiferromagnetic to the ferromagnetic regime, while keeping $J_{z}=1$ and all other parameters fixed. For $J=0$, the spins are uncoupled beyond the $s_{z}S_{z}$ term. As $J$ becomes increasingly negative (ferromagnetic $XY$ coupling), negativity increases, reflecting the stronger singlet-like correlations favored by the transverse exchange. In this case as well, entanglement death is observed with rising temperature, indicating that Negativity is highly susceptible to thermal fluctuations. MIN follows a qualitatively similar trend but remains nonzero even for smaller values of $J$ at higher $T$, underscoring its sensitivity to nonclassical disturbances beyond strict entanglement. The high $T$ “tail” of MIN is notably broader. At $T=0$, UIN again starts from a higher baseline and shows no abrupt thresholds. Its monotonic behavior across the entire $T$-range demonstrates its robustness as a measure of quantum correlations derived from information-theoretic uncertainty. Even at low temperature, a minimum ferromagnetic strength of J=0.7J=-0.7 is not sufficient for the CHSH parameter to exceed the classical limit. With the variation of $J$, the fragility of the quantumness measures follows the same hierarchy as in the previous case, namely $\text{Bell nonlocality}\subseteq\text{Negativity}\subseteq\text{UIN}(\text{MIN})$. This demonstrates that Bell nonlocality is the most challenging to sustain, while MIN and UIN serve as more reliable indicators of nonclassical correlations even at finite temperatures.
@@
Comparing Figs. 3 and 4 shows that for the strong ferromagnetic transverse coupling (J=2.5J=-2.5) and negative longitudinal exchange ($J_{z}=-1$), tuning the qubit Zeeman term ($B_{1}$) produces larger low-temperature entanglement and a wider Bell-violation window than tuning the qutrit Zeeman term ($B_{2}$). In both scans, MIN and UIN remain appreciable at higher temperatures, where entanglement and Bell nonlocality have already been suppressed, with UIN being the most thermally robust information-theoretic signature. The qualitative fragility ordering Bell $\subseteq$ Negativity $\subseteq$ UIN (MIN) remains valid across both parameter sweeps. However, the exact field windows and temperature thresholds differ depending on whether the qubit or qutrit field is being varied.
@@ 5 Conclusions and outlook
$\text{Bell nonlocality}\subseteq\text{Negativity}\subseteq\text{UIN}(\text{MIN}),$